OpenStudy (anonymous):

Nathan had an infection, and his doctor wanted him to take penicillin. Because Nathan’s father and paternal grandfather were allergic to penicillin, Nathan has a 75% chance of having the same allergy. The doctor performed a skin test to see whether Nathan would react to it. The test is 98% accurate. What is the probability that Nathan is not allergic to penicillin and the test predicts it?

OpenStudy (anonymous):

Answer choices: 0.005 0.015 0.245 0.735

OpenStudy (anonymous):

@jim_thompson5910 @SithsAndGiggles

OpenStudy (anonymous):

@jim_thompson5910 Can you please find someone?

jimthompson5910 (jim_thompson5910):

I found this page http://en.wikipedia.org/wiki/Accuracy_and_precision and I'm reading up on it. It might be able to lead us to the answer

OpenStudy (anonymous):

@jim_thompson5910 im really confused

jimthompson5910 (jim_thompson5910):

now I found this page and I think I'm onto something, so one more second

jimthompson5910 (jim_thompson5910):

oops sry this page http://vassarstats.net/bayes.html

OpenStudy (anonymous):

@jim_thompson5910 do you know what to do?

jimthompson5910 (jim_thompson5910):

It seems like something is missing here. Are you able to post a screenshot of this problem?

jimthompson5910 (jim_thompson5910):

I got this far, but I'm missing P(R|~A) A = Nathan is allergic R = test predicts Nathan has allergies P(A) = 0.75 P(~A) = 0.25 P(R|A) = 0.98 P(~R|A) = 0.02 P(R) = P(R|A)*P(A) + P(R|~A)*P(~A) P(R) = 0.98*0.75 + P(R|~A)*0.25

jimthompson5910 (jim_thompson5910):

Ok I got this A = Nathan is allergic T = test is positive P(A) = 0.75 (given) P(~A) = 0.25 (1-0.75) P(T) = 0.98 P(T|A) = 0.98 P(T|~A) = 0.02 we want to find P(~A|T) which is basically the probability of not having the allergy given the test is positive P(~A|T) = ( P(T|~A) * P(~A) )/( P(T) ) P(~A|T) = ( 0.02 * 0.25 )/( 0.98 ) P(~A|T) = 0.00510204081632 P(~A|T) = 0.005

jimthompson5910 (jim_thompson5910):

so the answer may be 0.005

OpenStudy (anonymous):

i just tried that. thats not the answer @jim_thompson5910

jimthompson5910 (jim_thompson5910):

then I'm not sure, I'll give it more thought kropot72 might be able to help though

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

 The probability that Nathan is not allergic to penicillin and the test predicts it is: \[\large 0.25\times0.02=0.005\]

OpenStudy (kropot72):

However you can also interpret the question as asking for the probability that the test correctly predicts that Nathan is not allergic. In this case the probability is: \[\large 0.25\times0.98=0.245\]

OpenStudy (kropot72):

I am sure that my second interpretation is correct: Required probability is 0.245.

OpenStudy (anonymous):

Yes you are right. It is 0.245. Thank you!